# Network Intrusion Datasets: A Survey, Limitations, and Recommendations

### v2

@@ 6.1 Choosing a Dataset
While many studies have highlighted NID data limitations, Flood et al. [8] introduced a practical framework for auditing network data. Given network flows, their approach proposes various metrics to detect mislabeling, artifacts, and poor data diversity. In addition to automated tests, they also conduct several manual checks. Replicating these steps can help identify a dataset’s drawbacks and assess its suitability.
@@ 6.2 Creating a Dataset
Correctness [32, 11] 1 – 10 Error-free implementation of the dataset creation steps – e.g., no packet losses, correct feature extraction and labeling, documentation corresponds to the data.
Diversity [33, 14, 105] 2, 3 Diversity and variability of both malicious traffic (multiple launches of the same attack with differing configurations) and benign traffic (rich usage profiles).
Documentation [45, 14, 105] 9 Documentation (metadata) provides detailed and unambiguous information about the dataset.
Extensibility [207] 2, 3, 9, 10 New scenarios, attacks, and traffic capture events can be added on demand.
Label completeness [23, 14, 105, 11] 6 All records are labeled directly, correctly, and unambiguously.
Pre-split [14] 2 – 4 Provides a predefined train-test split to support fair and objective benchmarking.
Privacy preservation [147, 14, 32] 2 – 5, 8 The data does not compromise the privacy (reveal sensitive information) of users. Captures from real networks are properly anonymized.
Realism [17, 144, 32, 54] 1 – 5, 7 The dataset contents and their characteristics reflect real-world networks. The environment is properly and completely configured. The data does not exhibit any unintended properties (e.g., artifacts) as a result of simulation.
Relevance [45, 14, 110] 1 – 6 Dataset scenarios (e.g., network topology, attacks, normal behavior) and the data itself (format, features, labels) are appropriate and relevant for the specified dataset objectives.
Reproducibility [37, 45, 208] 2, 3, 9, 10 The dataset contents can be regenerated on demand.
Timeliness [17, 14, 11] 1 – 3, 10 Contains up-to-date modern attacks and traffic profiles resembling the current threat landscape and the behavior of contemporary computer networks.
@@ 6.2.1 Preparation
    Further considerations should be given to diversity, ensuring the dataset includes diverse and variable normal and malicious traffic profiles. Insufficient variability can make the distinction between normal and malicious traffic trivial, leading to unrealistically overconfident NIDS performance. Extending capture duration by covering day-and-night or workweek-weekend cycles, such as in UGR'16 [105], is one way to enhance traffic variability. Although emulation and synthetic traffic generation methods are becoming more sophisticated, some argue that real background traffic needs to be used to achieve realism [14, 105]. However, it presents challenges like privacy preservation and establishing reliable ground truth [45].
@@ 6.2.2 Data Collection
    Execution & collection: After setting up the environment, the scenarios are run according to the execution plan, and relevant traffic is captured. The data format depends on the dataset goals, but multiple data formats (and sources), such as packets, flows, and logs, are generally preferred to support benchmarking of various NIDS types [144]. However, collecting data from multiple sources presents challenges, such as clock synchronization and the need for event correlation.
@@ 6.2.3 Post-Processing
    Data validation: Validating the data ensures correctness and realism – key factors in quality assurance. Data correctness refers to error-free capture and alignment with the designed scenarios. This involves investigating for packet losses or malformations, verifying proper attack execution and its desired effects (e.g., a DoS attack causing a service disruption), correct feature extraction (e.g., no unexpected N/A values, empty columns), the post-processing not introducing any artifacts or biases, and correct labeling. These checks are typically performed manually based on the data’s nature and semantics, although automated tests, as proposed by Flood et al. [8], are also viable. In addition to verifying data correctness, their realism can be validated using statistical tests along with domain-knowledge checks, as presented by Schoen et al. [56].
@@
    Data cleaning: In NID data creation, cleaning primarily refers to privacy preservation – i.e., anonymizing IP addresses and packet payloads to address ethical, privacy, and legal concerns for data collected from real-world networks. Additional information, like DNS queries, timestamps, or transport layer ports, may also need to be stripped. In some cases, removing entire erroneous or undesirable entries might even be necessary. However, excessive anonymization can reduce realism, limiting the dataset’s value [207]. For this reason, we advocate for minimal changes to avoid the degradation of realism.
@@ 6.2.4 Publication
    Comprehensive documentation elaborating on the above points would clarify the dataset’s properties, use cases, and limitations. Despite strict space constraints in scientific papers, we encourage researchers to provide this information as additional supplementary material. Although not perfect, AWID3 [139] serves as an example of decent documentation.
@@
    Sharing: The final step in dataset creation is sharing its data and metadata. We recommend uploading it to at least two (ideally multiple) data storage locations, including at least one public repository (Section 5.3.1) for redundancy. In addition, publishing source code and configurations of tools for data creation (i.e., environment setup, traffic generation and collection, feature extraction, and labeling) is also highly desirable to facilitate data reproducibility and validation [208]. If possible, both raw packets and extracted features should be shared.
@@ 6.3 Using a Dataset
You cannot see the future: Given the dynamic nature of network traffic (Section 3.1), benign and malicious traffic patterns can change over time. For this reason, it is crucial to respect the temporal properties of the data during evaluation – training data timestamps should strictly precede evaluation data to avoid temporal data snooping. Evaluation methods ignoring this aspect (including cross-validation) might introduce bias, leading to over-optimistic results [15, 39].
@@
The heaviest hammer is not always the best: As shown by numerous studies [23, 215, 216], even simplistic tree-based models like Random Forest or Gradient Boosting achieve near state-of-the-art performance with significantly faster execution. D’Hooge et al. demonstrated that even single-rule supervised ML models perform well on most cybersecurity datasets [217] due to the over-correlation of features with labels [12]. Therefore, a general recommendation is to start with simple models before employing deep learning [33]. For practical real-time intrusion detection, it is desirable to keep models simple so their inference (i.e., attack detection) is quick, enabling efficient processing of large traffic volumes. Therefore, we suggest adopting a data-centric perspective [218], focusing on understanding and enhancing the data with simple, explainable models rather than maximizing a specific metric with complex black-box models without understanding the data.
All that glitters is not gold: As discussed in the previous paragraph, even simplistic detection models can achieve decent performance, such as accuracy above $0.99$. However, this does not imply their actual usefulness. Consider an example of imbalanced traffic with a malicious-to-benign ratio of 1:100. Even if we classify all samples as benign, we receive $0.99$ accuracy while not detecting any attack. On the other hand, if an attack detection rate was 100% with only 1% of false alarms, every second alarm produced by the model would be incorrect. Due to this phenomenon, known as the base rate fallacy [22, 18], and the typical NID traffic imbalance, selecting the right metrics and their correct interpretation is crucial to gauge the overall benefit of the proposed system.
@@
Tools for network traffic replay like tcpreplay [222] preserve packet contents while unifying time dependencies within a single environment, thus reducing time-related biases. However, merging distinct datasets can still introduce artifacts. For instance, data from different systems (e.g., Linux vs. Windows) might present differences in packet contents, such as the employed encryption protocols. Additionally, replaying complex attacks requiring both sides of the communication might be challenging, as most replay tools are stateless [185]. These factors can introduce additional hidden biases during data merging. For this reason, NIDS methods should also be benchmarked on other datasets originating from a single source to prevent potential biases from skewing the results.
@@ 7.1 Data Generation
In this matter, a crucial open problem to address is how to enable continual real-world network data capture and annotation while preserving privacy and ensuring label accuracy. Previous works such as Kyoto 2006+ [99], MAWILab [50], and LuFlow [148] have explored this approach but are no longer updated. Currently, we are not aware of any similar project for continuous data capture providing labeled traces for NID purposes. For this reason, future research should explore automated, privacy-preserving data collection methods from real-world environments with accurate labeling.
@@
In addition to previous scenarios, Hindy et al. [46] and Keersmaeker et al. [229] pointed out that many attacks and IoT environments are not covered within public NID datasets at all. For these reasons, future research should prioritize covering specific scenarios and environments, along with ensuring robust coverage of emerging threats to streamline NIDS research in these areas. However, without continually captured and extensible datasets is made, new datasets will remain susceptible to getting outdated due to the ever-changing nature of NID.
@@ 7.2 Data Processing
This strong focus on flow-based data has led to a relative neglect of research on alternative data granularities, such as packet-level features, time/packet window statistics, and cross-packet contents. These could help detect attacks missed by flow-based methods [79]. Future research should thus explore these alternatives and their fusion to improve detection performance, robustness, and adaptability to evolving threats.
@@ 7.3 Data Validation
A recent paper by Flood et al. [8] proposed several manual and automated tests to audit common limitations of NID data, such as mislabeling, artifacts, and lack of diversity. Although this represents a significant step toward data quality evaluation and the selection of suitable datasets, their heuristics focus only on a subset of typical problems, while the proposed manual analysis requires access to both flow-based and raw data.
In order to improve NID data evaluation, future research should formalize methods and develop frameworks for evaluating NID data validity and quality, i.e., ensure datasets are both realistic and error-free. Current data validation heavily relies on manual analysis, as seen in studies on CIC-IDS2017 and CSE-CIC-IDS2018 issues [108, 48, 165], with Engelen et al. [226] recommending manual analysis by domain experts before public data release. Such validations are time-consuming and dependent on the availability of experts, posing additional challenges for data creators. Therefore, research on NID data quality should also integrate expert knowledge into (semi-)automated validation methods, potentially building on the Flood et al.’s work [8]. Standardized approaches to verify other aspects of data quality – e.g., the validity of features via domain knowledge, documentation via Large Language Models (LLMs) [233], or the detection of common issues – would benefit data creators to validate their datasets and assist users in selecting suitable data, enhancing NID research relevance, trustworthiness, and explainability.
@@ 7.4 Data Publishing
One of the potential solutions is the development of a centralized indexing system for NID datasets, similar to the digital object identifier (DOI), as proposed by Abt et al. [10] in 2014 and later emphasized by Ring et al. [14]. Such a system would simplify data discovery, provide persistent availability, and reduce delivery delays. Despite these benefits and multiple calls, we are not aware of any indexing system, data storage, or publicly maintained repository of (network) intrusion datasets.
@@ 8 Related Work
This section compares the presented survey with other works discussing network intrusion detection datasets. While data-specific NID surveys (this paper) are rare, general NID surveys often include datasets as well. Naturally, they typically lack detailed discussion on data-specific topics yet still provide a relevant reference point for many readers. We thus compare our work to those focused explicitly on NID data and five general surveys with data-related sections, summarized in Table 5. Based on the related work analysis, we conclude that our paper:
@@
The closest study to ours, and its inspiration in many aspects, is the work by Ring et al. [14]. Considered one of the first surveys on NID data, it formalized data properties, compared existing datasets, and provided data-related recommendations. However, since it was published in 2019, it does not include newer datasets. Our study builds on this work by refining some established properties (e.g., removing the data balance property, as NID data are naturally imbalanced) and incorporating newly released datasets. We also introduce sections on data limitations and trends to address recent findings and emphasize data quality, aiming to provide a holistic view of NID data.
@@
Although Hindy et al. [46] cover aspects similar to this paper, they primarily focus on the network threat landscape and its relation to 21 NID datasets (until 2018). The study further discusses NIDS algorithms’ and datasets’ popularity, threat taxonomy, IDS limitations, and recommendations for future research. In contrast, our study includes more datasets, compares them by their properties rather than by attacks, and provides a more detailed discussion of NID data-related aspects.
@@
As mentioned in Section 2.1, our survey focuses on network intrusion datasets regardless of the environment, with the key condition being that the data must be captured from the network. Therefore, we exclude IDS datasets where most of the data originates from non-network sources like end-host devices or sensors, as typical for Industrial Control Systems (ICS) or Internet of Things (IoT) datasets. Although relevant for environment-specific intrusion detection, we consider such datasets out-of-scope. For this reason, we recommend exploring other, more specialized surveys for data in specific areas. For instance, Conti et al. [236] and Koay et al. [237] review ICS testbeds and datasets, whereas IoT-specific datasets are covered by Keersmaeker et al. [229] and Kaur et al. [238].
@@ 9 Conclusions
In total, this survey has surveyed 89 NID datasets, extracting 13 properties via manual analysis of their documentation and data. The resulting comparative table is a valuable resource to aid in data selection for benchmarking new NIDS proposals. Analyzing these properties, we revealed several data creation trends, such as the growing number of published datasets each year, a preference for emulation for data generation, and the rising popularity of CICFlowMeter features.
@@ A.1 Data Search and Filtering Process
    Hence, we work with a mixture of scientific papers and other resources that do not follow a standard scientific paper format or provide experimental results, making common research quality assessments [239] impractical. Since we do not aim to propose new NIDS data quality metrics, our quality check is based on a simple premise: Is the dataset useful for potential users? We thus specify:

### v3

@@ 6.1 Choosing a Dataset
While many studies merely point out limitations of NID data, Flood et al. [73] present a practical framework for auditing network datasets. Given network flows, their approach proposes various metrics to detect mislabeling, artifacts, and poor data diversity. Alongside automated tests, they also conduct several manual checks. Replicating these steps can help identify a dataset’s drawbacks and assess its suitability.
@@ 6.2 Creating a Dataset
Correctness Viegas et al. [222] Małowidzki et al. [134] 1–10 Error-free implementation of the dataset creation steps – e.g., no packet losses, correct feature extraction and labeling, documentation corresponds to the data.
Diversity Catillo et al. [38] Ring et al. [178] Maciá-Fernández et al. [131] 2, 3 Diversity and variability of both malicious traffic (multiple launches of the same attack with differing configurations) and benign traffic (rich usage profiles).
Documentation Landauer et al. [115] Ring et al. [178] Maciá-Fernández et al. [131] 9 Documentation (metadata) provides detailed and unambiguous information about the dataset.
Extensibility Kenyon et al. [104] 2, 3, 9, 10 New scenarios, attacks, and traffic capture events can be added on demand.
Label completeness Apruzzese et al. [18] Ring et al. [178] Maciá-Fernández et al. [131] Małowidzki et al. [134] 6 All records are labeled directly, correctly, and unambiguously.
Pre-split Ring et al. [178] 2–4 Provides a predefined train-test split to support fair and objective benchmarking.
Privacy preservation Ferriyan et al. [71] Ring et al. [178] Viegas et al. [222] 2–5, 8 The data does not compromise the privacy (reveal sensitive information) of users. Captures from real networks are properly anonymized.
Realism Apruzzese et al. [16] Myneni et al. [152] Viegas et al. [222] Shiravi et al. [195] 1–5, 7 The dataset contents and their characteristics reflect real-world networks. The environment is properly and completely configured. The data does not exhibit any unintended properties (e.g., artifacts) as a result of simulation.
Relevance Landauer et al. [115] Ring et al. [178] Ring et al. [176] 1–6 Dataset scenarios (e.g., network topology, attacks, normal behavior) and the data itself (format, features, labels) are appropriate and relevant for the specified dataset objectives.
Reproducibility Catillo et al. [37] Landauer et al. [115] Lewandowski [123] 2, 3, 9, 10 The dataset contents can be regenerated on demand.
Timeliness Apruzzese et al. [16] Ring et al. [178] Małowidzki et al. [134] 1–3, 10 Contains up-to-date modern attacks and traffic profiles resembling the current threat landscape and the behavior of contemporary computer networks.
@@ 6.2.1 Preparation
    Further considerations should be given to diversity, ensuring the dataset includes diverse and variable normal and malicious traffic profiles. Insufficient variability can make the distinction between normal and malicious traffic trivial, leading to unrealistically inflated NIDS performance. Extending capture duration by covering day-night or workweek-weekend cycles, as in UGR'16 [131], is one way to enhance traffic variability. Although emulation and synthetic traffic generation methods are becoming more sophisticated, some argue that real background traffic must be used for realism [178, 131]. However, it presents challenges like privacy preservation and establishing reliable ground truth [115].
@@ 6.2.2 Data Collection
    Execution & collection: After setting up the environment, the scenarios are run according to the execution plan, and relevant traffic is captured. The data format depends on the dataset goals, but multiple data formats (and sources), such as packets, flows, and logs, are preferred to support benchmarking of various NIDS types [152]. However, collecting data from multiple sources presents challenges, such as clock synchronization and the need for event correlation.
@@ 6.2.3 Post-Processing
    Data validation: Validating the data ensures correctness and realism – key factors in quality assurance. Data correctness refers to error-free capture and alignment with the designed scenarios. This involves investigating for packet losses or malformations, verifying proper attack execution and its desired effects (e.g., a DoS attack causing a service disruption), correct feature extraction (e.g., no unexpected N/A values, empty columns), the post-processing not introducing any artifacts or biases, and correct labeling. These checks are typically performed manually based on the data’s nature and semantics, although automated tests, as proposed by Flood et al. [73], are also viable. In addition to verifying correctness, data realism can be validated using statistical tests along with domain-knowledge checks, as presented by Schoen et al. [188].
@@
    Data cleaning: In NID data creation, cleaning primarily refers to privacy preservation – i.e., anonymizing IP addresses and packet payloads to address ethical, privacy, and legal concerns for data collected from real-world networks. Additional information, like DNS queries, timestamps, or transport layer ports, may also need to be stripped. In some cases, removing entire erroneous entries might even be necessary. However, excessive anonymization can reduce realism, limiting the dataset’s value [104]. For this reason, we advocate for minimal changes to avoid the degradation of realism.
@@ 6.2.4 Publication
    Comprehensive documentation elaborating on the above points would clarify the dataset’s properties, use cases, and limitations. Despite strict space constraints in scientific papers, we encourage providing this information as additional supplementary material. Although not perfect, AWID3 [41] serves as an example of decent documentation.
@@
    Sharing: The final step in dataset creation is sharing its data and metadata. We recommend uploading them to multiple data storage locations, including at least one public repository (Section 5.3.1) for redundancy. In addition, publishing source code and configurations of tools for data creation (i.e., environment setup, traffic generation and collection, feature extraction, and labeling) is also highly desirable to facilitate data reproducibility and validation [123]. If possible, both raw packets and extracted features should be shared.
@@ 6.3 Using a Dataset
You cannot see the future: Given the dynamic nature of networks (Section 3.1), benign and malicious traffic patterns can change over time. Therefore, it is crucial to respect the temporal properties of the data during evaluation – training data timestamps strictly precede evaluation data to avoid temporal data snooping. Evaluation methods ignoring this aspect (including cross-validation) might introduce bias, leading to over-optimistic results [20, 167].
@@
The heaviest hammer is not always the best: As shown by numerous studies [18, 234, 223], even simplistic tree-based models like Random Forest or Gradient Boosting achieve near state-of-the-art performance with significantly faster execution. D’hooge et al. [63] demonstrated that even single-rule supervised ML models perform well on most cybersecurity datasets due to the over-correlation of features with labels [199]. Therefore, a general recommendation is to start with simple models before employing deep learning [38]. For practical real-time intrusion detection, it is desirable to keep models simple so their inference (i.e., attack detection) is quick, enabling efficient processing of large traffic volumes. Therefore, we suggest adopting a data-centric perspective [233], focusing on understanding and enhancing the data with simple, explainable models rather than maximizing a specific metric with complex black-box models without understanding the data.
All that glitters is not gold: As discussed in the previous paragraph, even simplistic detection models can achieve decent performance, such as accuracy above $0.99$. However, this does not imply their actual usefulness. Consider an example of imbalanced traffic with a malicious-to-benign ratio of 1:100. Even if we classify all samples as benign, we receive $0.99$ accuracy while not detecting any attack. On the other hand, if a detection rate was 100% with only 1% of false alarms, every second alarm produced by the model would be incorrect. Due to this phenomenon, known as the base rate fallacy [21, 9], and the typical NID traffic imbalance, selecting the right metrics and their correct interpretation is crucial to gauge the overall benefit of the proposed system.
@@
Tools for network traffic replay like tcpreplay [212] preserve packet contents while unifying time dependencies within a single environment, thus reducing time-related biases. However, merging distinct datasets can still introduce artifacts. For instance, data from different systems (e.g., Linux vs. Windows) might present differences in packet contents, such as in encryption protocols. Additionally, replaying complex attacks requiring both sides of the communication might be challenging, as most replay tools are stateless [3]. These factors can introduce additional hidden biases during data merging. For this reason, NIDS methods should also be benchmarked on other datasets originating from a single source to prevent potential biases from skewing the results.
@@ 7.1 Data Generation
In this matter, a key open problem is how to enable continual real-world network data capture and annotation while preserving privacy and ensuring label accuracy. Previous works such as Kyoto 2006+ [202], MAWILab [74], and LuFlow [142] have explored this approach but are no longer updated. Currently, we are not aware of any similar project for continuous data capture providing labeled traces for NID purposes. For this reason, future research should explore automated, privacy-preserving data collection methods from real-world environments with accurate labeling.
@@
In addition to previous scenarios, Hindy et al. [91] and De Keersmaeker et al. [54] pointed out that many attacks and IoT environments are not covered within public NID datasets at all. Therefore, future research should prioritize covering specific scenarios and environments, along with ensuring robust coverage of emerging threats to streamline NIDS research in these areas. However, without continually captured and extensible datasets is made, new datasets will remain susceptible to getting outdated due to the ever-changing nature of NID.
@@ 7.2 Data Processing
The focus on flow-based data has led to a relative neglect of research on alternative data granularities, such as packet-level features, time/packet window statistics, and cross-packet contents. These could help detect attacks missed by flow-based methods [220]. Future research should thus explore these alternatives and their fusion to improve detection performance, robustness, and adaptability to evolving threats.
@@ 7.3 Data Validation
A recent paper by Flood et al. [73] proposed manual and automated tests to audit common limitations of NID data, such as mislabeling, artifacts, and lack of diversity. Although this represents a significant step toward data quality evaluation and the selection of suitable datasets, their heuristics focus only on a subset of typical problems, while the proposed manual analysis requires access to both flow-based and raw data.
In order to improve NID data evaluation, future research should formalize methods and develop frameworks for evaluating NID data validity and quality, i.e., ensure datasets are both realistic and error-free. Current data validation heavily relies on manual analysis, as seen in studies on CIC-IDS2017 and CSE-CIC-IDS2018 issues [66, 127, 117], with Engelen et al. [65] recommending manual analysis by domain experts before public data release. Such validations are time-consuming and dependent on the availability of experts, posing additional challenges for data creators. Therefore, research on NID data quality should also integrate expert knowledge into (semi-)automated validation methods, potentially building on the Flood et al. [73]’s work. Standardized approaches to verify other aspects of data quality – e.g., the validity of features via domain knowledge, documentation via Large Language Models (LLMs) [81], or the detection of common issues – would benefit data creators to validate their datasets and assist users in selecting suitable data, enhancing NID research relevance, trustworthiness, and explainability.
@@ 7.4 Data Publishing
One of the potential solutions is the development of a centralized indexing system for NID datasets, similar to the digital object identifier (DOI), as proposed by Abt and Baier [2] and later emphasized by Ring et al. [178]. Such a system would simplify data discovery, provide persistent availability, and reduce delivery delays. Despite these benefits and multiple calls, we are not aware of any indexing system, data storage, or publicly maintained repository of (network) intrusion datasets.
@@ 8 Related Work
This section compares the presented survey with other studies discussing network intrusion detection datasets. While data-specific NID surveys (this paper) are rare, general NID surveys often mention datasets as well. Naturally, they typically lack detailed discussion on data-specific topics yet still provide a relevant reference point for many readers. We thus compare our work to those focused explicitly on NID data and five general surveys with data-related sections, summarized in Table 5. Based on the related work analysis, we conclude that our paper:
@@
The closest study, and our inspiration in many aspects, is Ring et al. [178]’s work. This foundational work on NID data formalized data properties, compared existing datasets, and provided data-related recommendations. However, it does not include newer datasets. Our study builds on it by refining some of its properties (e.g., removing the data balance property, as NID data are naturally imbalanced) and incorporating newly released datasets. We also introduce sections on data limitations and trends to address recent findings and emphasize data quality, aiming to provide a holistic view of NID data.
@@
Although Hindy et al. [91] cover aspects similar to this paper, they mainly focus on the network threat landscape and its relation to 21 NID datasets (until 2018). The study further discusses NIDS algorithms’ and datasets’ popularity, threat taxonomy, IDS limitations, and recommendations for future research. In contrast, our study includes more datasets, compares them by their properties rather than by attacks, and provides a more detailed discussion of NID data-related aspects.
@@
As mentioned in Section 2.1, this survey focuses on network intrusion datasets regardless of the environment, with the condition that the data must be captured from the network. Therefore, we exclude IDS datasets where most of the data originates from non-network sources like end-host devices or sensors, as typical for Industrial Control Systems (ICS) or Internet of Things (IoT) datasets. Although relevant for environment-specific intrusion detection, we consider such datasets out-of-scope. For this reason, we recommend exploring other, more specialized surveys for data in specific areas. For instance, Conti et al. [45] and Koay et al. [106] review ICS testbeds and datasets, whereas IoT-specific datasets are covered by De Keersmaeker et al. [54] and Kaur et al. [102].
@@ 9 Conclusions
In total, this paper has surveyed 89 NID datasets, extracting 13 properties via manual analysis of their documentation and data. The resulting comparative table is a valuable resource to aid in data selection for benchmarking new NIDS proposals. Analyzing these properties, we revealed several data creation trends, such as the growing number of published datasets each year, a preference for emulation for data generation, and the rising popularity of CICFlowMeter features.
@@ A.1 Data Search and Filtering Process
    Hence, we work with a mixture of scientific papers and other resources that do not follow a standard scientific paper format or provide experimental results, making common research quality assessments [Yang et al., 2021] impractical. Since we do not aim to propose new NIDS data quality metrics, our quality check is based on a simple premise: Is the dataset useful for potential users? We thus specify: